# TASAR: Transfer-based Attack on Skeletal Action Recognition

### v1

@@ A Post-train Bayesian Perspective on Attack
HDM05 ST-GCN No MG 96.81 13.23 58.58 38.97 17.89
MG 97.30 14.26 60.54 37.50 23.43
MS-G3D No MG 46.28 38.74 97.79 36.27 28.99
MG 44.89 47.62 100.00 36.49 36.34
NTU60 ST-GCN No MG 98.77 33.14 54.97 17.87 37.50
MG 99.29 42.45 64.60 20.33 49.41
MS-G3D No MG 40.60 48.69 99.61 37.51 46.07
MG 48.87 51.18 99.61 41.49 40.14
NTU120 ST-GCN No MG 98.53 20.83 24.46 14.71 21.81
MG 99.26 19.60 19.37 15.28 22.79
MS-G3D No MG 34.64 36.27 100.00 25.49 32.03
MG 41.16 47.28 100.00 28.83 40.60
Table 5: The ablation experiments of motion gradient. ‘MG’ means TASAR with motion gradient, ‘No MG’ means TASAR without using motion gradient.
Surrogate Transferability
It is widely believed that transfer-based attacks in S-HAR are highly sensitive to the surrogate choice (Lu et al. 2023; Wang et al. 2023, 2021). In this subsection, we provide a detailed analysis of the factors contributing to this phenomenon. When looking at the results in Tab. 1 and the visualization of loss landscape in Fig. 2 and Appendix B, we note that loss surface smoothness correlates with the adversarial transferability. For example, CTR-GCN, manifesting smoother regions within the loss landscape, demonstrates higher transferability than ST-GCN and STTFormer. STTFormer trained on NTU 120 has a smoother loss surface than ST-GCN (see Appendix B), resulting in higher transferability than ST-GCN. For NTU 60, STTFormer shows a similar loss surface to that of ST-GCN and exhibits comparable transferability.Therefore, we suspect that the loss surface smoothness plays a pivotal role in boosting adversarial transferability for S-HAR, potentially outweighing the significance of gradient-based optimization techniques. Next, two-stream MS-G3D shows the highest transferability. Unlike other surrogates, which solely extract joint information, MS-G3D uses a two-stream ensemble incorporating both joint and bone features, thereby effectively capturing relative joint movements. In conclusion, we suggest that skeletal transfer-based attacks employ smoother two-stream surrogates incorporating both joint and bone information.
Conclusion

### v5

@@ C.2 Additional Results
0.001 99.26 19.60 19.37 15.28 22.79 63.60
0.01 95.88 17.62 13.10 14.29 20.23 64.30
0.05 96.43 17.26 13.10 14.88 21.43 63.60
0.1 96.43 17.62 13.10 13.69 21.43 60.56
1 96.43 18.45 11.90 16.07 23.21 60.39
1.5 96.43 18.45 13.10 15.48 23.81 56.58
2 97.02 17.26 13.10 16.67 22.02 54.61
Figure 9: The success rate with different perturbation size. The surrogate model is uniformly chosen as MSG-3D and the dataset is NTU60.
Perturbation Budget
In this section, we analyze the impact of attack strength on adversarial transferability. We increase the perturbation budget from 0.01 to 0.05, the results are shown in Figure 9. The general pattern of the lines aligns with the threshold setting, indicating that a larger perturbation budget consistently enhances adversarial transferability across various surrogate models.
The Visual Quality of Targeted Adversarial Examples
TASAR can successfully attack the original class to a target with an obvious semantic gap without being detected by humans. We show an additional visual example in Figure 8.
Comparsion with Different Training Strategies.
We conducted additional experiments to compare the performance of modeling $p\left(\theta^{\prime}\mid D,\theta\right)$ , $p(\theta\mid D)$ and $p\left(\theta,\theta^{\prime}\mid D\right)$. The default setting of TASAR corresponds to $p\left(\theta^{\prime}\mid D,\theta\right)$, where $\theta^{\prime}$ is trainable while $\theta$ remains fixed. $p(\theta\mid D)$ represents a standard Bayesian neural network without adding appended models. For this case, we train the Bayesian surrogates with a cyclical variant of Stochastic Gradient Markov Chain Monte Carlo[1] to sample 3 models from the posterior distribution of neural network weights. For $p\left(\theta,\theta^{\prime}\mid D\right)$, 9 appended models are added behind BNNs and both $\theta$ and $\theta^{\prime}$ are trainable. Due to the high optimization complexity(updating both $\theta$ and $\theta^{\prime}$), we use vanilla post-train Bayeian optimization instead of improved post-train dual Bayesian optimization optimization for updating $\theta^{\prime}$.
Table 6: The attack success rate(%) of untargeted attacks on NTU60.
Surrogate Method NTU60
ST-GCN 2s-AGCN MS-G3D CTR-GCN FR-HEAD SFormer
ST-GCN $p(\theta^{\prime}|D,\theta)$ (TASAR) 99.29 42.55 64.60 20.33 49.41 17.22
$p(\theta|D)$ 90.63 37.11 69.53 17.58 43.36 25.39
$p(\theta,\theta^{\prime}|D)$ 93.75 41.80 73.04 18.75 46.87 22.65
Table 7: Model Size and Training Time measurement on NTU60. ’MS’ means The Model Size(M) and ’TT’ means The Training Time(hours).
Surrogate Method NTU60
MS TT
ST-GCN $p(\theta^{\prime}|D,\theta)$ 3.54 0.65
$p(\theta|D)$ 9.30 5.7
$p(\theta,\theta^{\prime}|D)$ 9.36 6.4
The results are presented in Table 6. Compared to modeling $p(\theta\mid D)$ and $p(\theta,\theta^{\prime}|D)$, TASAR achieves superior performance in transfer attacks on three out of five black-box models and demonstrates the best white-box attack performance on ST-GCN. This advantage arises from the use of improved post-training Dual Bayesian optimization, which enables smoothed posterior sampling and improves adversarial transferability. Moreover, unlike modeling $p(\theta\mid D)$ and $p(\theta,\theta^{\prime}|D)$, where averaging gradients from multiple surrogates diminish white-box attack effectiveness, our post-training strategy preserves the pre-trained model intact, not reducing the original white-box attack strength. Further, TASAR significantly reducing computational overhead and accelerating training process, as shown in Table 7.
Quantification of Model Smoothness.
We quantitatively measure the changes in loss as $\mathbf{x}$ is perturbed along a random direction with varying magnitudes. Specifically, we first sample $\mathbf{d}$ from a Gaussian distribution and normalize it onto the $\ell_{2}$ unit norm ball as $\mathbf{d}\leftarrow\frac{\mathbf{d}}{\|\mathbf{d}\|_{F}}$. Then, we calculate the loss change (smoothness) $f(a)$ for different magnitudes $a$:
$f(a)=\left|L(\mathbf{x}+a\cdot\mathbf{d},y,\theta)-L(\mathbf{x},y,\theta)\right|.$ (21)
Table 8: Loss changes ($f(a)$) measurement for normally trained Surrogate and TASAR on HDM05. ”NT” means ”Normally Training”.
Surrogate Method Magnitude
-1.0 -0.8 -0.6 -0.4 0.4 0.6 0.8 1.0
ST-GCN NT 7.46 6.05 4.34 2.08 2.06 4.33 5.91 7.24
TASAR 2.66 1.70 1.03 1.77 2.27 1.40 1.65 2.31
We calculate $f(a)$ 20 times with different randomly sampled $\mathbf{d}$, and take the average results. For fair comparison, we use the same $\mathbf{d}$ each sampling in both ‘NT’ and ‘TASAR’. The experimental results reveal that TASAR achieves a significantly smoother loss landscape compared to Normally Train (NT) across all magnitudes of perturbation. For large perturbations ($|a|=1.0$), TASAR reduces the loss change by approximately threefold compared to NT. Additionally, TASAR maintains a more uniform smoothness across different magnitudes, while NT exhibits sharper variations, with larger loss changes. This indicates that TASAR effectively smoothens the loss landscape, contributing to improved transferability.
Appendix D Detailed RobustBenchHAR Settings
Here we report the detailed experimental settings in our experiments. All experiments are conducted on one NVIDIA GeForce RTX 3090 GPU.
(A) Datasets.
We choose three widely adopted datasets: NTU60 (Shahroudy et al., 2016) , NTU120 (Liu et al., 2019) and HDM05(Müller et al., 2007). The HDM05 dataset comprises 130 action classes and includes 2337 sequences performed by 5 non-professional actors. NTU60 offers 60 action classes, it comprises 56,880 videos captured from 40 subjects across 155 camera viewpoints. NTU120 extends NTU60 with 120 action classes, it contains 114,480 videos from 106 subjects across 155 camera viewpoints. Due to variations in data pre-processing settings among S-HAR classifiers (such as data requiring subsampling(Zhang et al., 2019b)), we unify the data format following Wang et al. (2023). For NTU60 and NTU120, we subsample frames to 60. For HDM05, we segment the data into 60-frame samples(Diao et al., 2021).
(B) Evaluated Models.
We evaluate TASAR in three categories of surrogate/victim models.
(1) Normally trained models: We adapt 5 commonly used GCN-based models, i.e., ST-GCN(Yan et al., 2018), 2S-AGCN(Shi et al., 2019a), CTR-GCN(Chen et al., 2021), MS-G3D(Liu et al., 2020b), FR-HEAD(Zhou et al., 2023) and 2 Transformer-based models, i.e., STTFormer(Qiu et al., 2022) and SkateFormer(Do & Kim, 2024). Below we introduce the seven skeletal classifiers in details. ST-GCN (Yan et al., 2018) is the first time to apply graph convolution to S-HAR, constructing nodes and edges in the topology using skeletal information. CTR-GCN (Chen et al., 2021) improves the design of GCNs of ST-GCN and proposes to dynamically learn different topologies by learning a shared topology as a common prior for all channels and refining it for each channel. 2s-AGCN (Shi et al., 2019a) enables the model to learn graph topologies end-to-end through self-attention. It also incorporates a dual-stream framework to model first-order and second-order information simultaneously. MS-G3D (Liu et al., 2020b) proposes a disentangled multi-scale aggregation scheme to eliminate redundant dependencies between node features from different neighborhoods, thereby capturing global joint relationships on human skeletons. FR-HEAD (Zhou et al., 2023) applies contrastive feature refinement at various stages of GCNs to build multi-level feature extraction. This allows ambiguous samples to be dynamically discovered and calibrated in the feature space. STTFormer (Qiu et al., 2022) divides the skeleton sequence into temporal tuples to capture the relationships between different joints in consecutive and non-adjacent frames. SkateFormer (Do & Kim, 2024) classifies essential skeletal-temporal relationships for S-HAR into four distinct categories and utilizes self-attention in each partition to focus on key joints and frames crucial for recognition. To the best of our knowledge, this is the first work to investigate the robustness of Transformer-based S-HARs.
(2) Ensemble model: An ensemble of ST-CGN, MS-G3D and DGNN (Shi et al., 2019b).
(3) Defense models: We employ BEAT (Wang et al., 2023) and TRADES (Zhang et al., 2019a), which all demonstrate their robustness for skeletal classifiers. BEAT (Wang et al., 2023) proposes a black-box defense framework, which transforms any pre-trained classifier into a more robust one by maximizing the joint probability of clean data, adversarial examples and the classifier through joint Bayesian treatments.TRADES (Zhang et al., 2019a) is a white-box defense method that introduces a KL-divergence loss function for adversarial training. TRADES not only accounts for natural error but also incorporates adversarial error, balancing robustness and accuracy.
(C) Baselines.
Unlike images or videos, the space available for attacking skeletons is much smaller, making adversarial perturbations on skeletons more easily detectable, here we choose two state-of-the-art attacks against S-HAR: (1) SMART (Wang et al., 2021) ensures the imperceptibility of the attack by introducing an adversarial attack perceptual loss function for S-HAR. (2) CIASA (Liu et al., 2020a)maintains the spatio-temporal constraints of joint connections and skeletal bone lengths through spatial skeleton realignment and further ensures the anthropomorphic plausibility by utilizing GAN(Goodfellow et al., 2014a) for regularization.
Besides the attacks specifically designed for S-HAR, we also select general transfer-based attacks as baselines, these attacks include (1) Gradient-based attacks, such as I-FGSM (Kurakin et al., 2018), an iterative fast gradient method; MIFGSM (Dong et al., 2018), which integrates momentum into I-FGSM to stabilize the update direction and prevent getting stuck in poor local maxima; and the latest MIG (Ma et al., 2023), which uses integrated gradients to steer the generation of adversarial perturbations and adjusts them according to the momentum updating strategy. (2) Input transformation attacks, such as DIM (Xie et al., 2019), which improves the transferability of adversarial examples by creating diverse input patterns. (3) Ensemble-based/Bayesian attacks, including ENS (Dong et al., 2018), which attacks multiple models with fused logit activations; SVRE (Xiong et al., 2022), which escapes poor local optima by computing an unbiased gradient estimate through variance reduction for each iteration; and BA (Li et al., 2023), which fine-tunes the weights of the surrogate model in a Bayesian manner, thereby creating an ensemble of infinitely many DNNs as surrogates.
For a fair comparison, we run 200 iterations for all attacks under $l_{\infty}$ norm-bounded perturbation of size 0.01. For TASAR, we use the iterative gradient attack instead of FGSM.
(D) Implementation Details Of TASAR.
Our appended model is a simple two-layer fully-connected layer network. Unless specified otherwise, we use $K=3$ and $M=20$ in TASAR for default and explain the reason in the ablation study later.
[…]
During the dual Bayesian optimization, we set $\gamma=0.1/\left\|\Delta\theta_{*}^{\prime}\right\|_{2}$ and perform training for 5 epochs with a learning rate of 0.03. Additionally, we always set $\lambda_{\varepsilon,\sigma}=0.001$.
During Inference, SWAG adjusts the covariance matrix using a constant multiplier to decouple the learning rate from covariance(Maddox et al., 2019). Here we always use 1.5 as the rescaling factor. When performing attacks, we set $\sigma=0.009$ for models with SWAG. The $w_{1}$,$w_{2}$ and $w_{3}$ are set as 0.8, 0.1 and 0.1.
(E) Computing Resource.
The experimental platform used in this study is equipped with an AMD EPYC 7542 32-Core CPU operating at a clock speed of 2039.813 GHz, four NVIDIA GeForce RTX 3090 GPUs, and 24 GB of memory per GPU. The proposed method was implemented using the open-source machine learning framework PyTorch.